An electrical contractor’s records during the last five weeks indicate the number of job requests: Week: 1 2 3 4 5 Requests: 20 22 18 21 22 Predict the number of requests for week 6 using each of these methods: a. Naive. b. A four-period moving average. c. Exponential smoothing with .30. Use 20 for week 2 forecast.
Answer #1

a)

Using Naive method, forecast for week 6 = Actual number of requests for previous week (week 5)

= 22

---------------------------------------

b)

Four-period moving average forecast for week 6 = Average of previous 4 weeks

= (22+18+21+22)/4

= 20.75

---------------------------------------

c)

Exponential smoothing forecast is calculated by the following formula:

Ft = *At-1 + (1-)*Ft-1

where, Ft is the forecast of period t, and At is the actual demand of period t,

Forecast for week 3 = 0.3*22+(1-0.3)*20 = 20.60

Forecast for week 4 = 0.3*18+(1-0.3)*20.60 = 19.82

Forecast for week 5 = 0.3*21+(1-0.3)*19.82 = 20.17

Forecast for week 6 = 0.3*22+(1-0.3)*20.17 = 20.72